# Controllable Collision Scenario Generation via Collision Pattern Prediction

### v1

@@ Abstract
Evaluating the safety of autonomous vehicles (AVs) requires diverse, safety-critical scenarios, with collisions being especially important yet rare and unsafe to collect in the real world. Therefore, the community has been focusing on generating safety-critical scenarios in simulation. However, controlling attributes such as collision type and time-to-accident (TTA) remains challenging. We introduce a new task called controllable collision scenario generation, where the goal is to produce trajectories that realize a user-specified collision type and TTA, to investigate the feasibility of automatically generating desired collision scenarios. To support this task, we present COLLIDE, a large-scale collision scenario dataset constructed by transforming real-world driving logs into diverse collisions, balanced across five representative collision types and different TTA intervals. We propose a framework that predicts Collision Pattern, a compact and interpretable representation that captures the spatial configuration of the ego and the adversarial vehicles at impact, before rolling out full adversarial trajectories. Experiments show that our approach outperforms strong baselines in both collision rate and controllability. Furthermore, generated scenarios consistently induce higher planner failure rates, revealing limitations of existing planners. We demonstrate that these scenarios fine-tune planners for robustness improvements, contributing to safer AV deployment in different collision scenarios.
@@ II Related Work
Scenario Generation. Existing approaches can be categorized as data-driven, adversarial, and knowledge-based, according to [1]. Data-driven methods [23, 24, 25] extract rare collision events from large-scale real-world datasets [26], but such events are sparse and highly imbalanced in terms of both collision type and urgency (TTA), limiting their value for training or evaluation. Adversarial methods [7, 27, 8] perturb agent trajectories within learned traffic dynamics models [28]. Although effective at generating failures, they are biased toward frequent outcomes (e.g., rear-end or head-on), since traffic priors and minimal-change objectives favor straight-line or opposing motions over rarer intrusions. Knowledge-based approaches [29, 30, 17, 31] incorporate rule-based priors or semantic graphs to constrain the generation process. For instance, SafeBench [31] and RiskBench [17] rely on pre-defined templates, which guarantee coverage of certain collision types but are not scalable, as increasing intra-class variation requires substantial manual effort and does not generalize efficiently. Similarly, CausalAF [29] is confined by a hand-crafted causal graph, preventing diverse variations or generalization to other scenarios.
@@ IV-B Framework Overview
As illustrated in Fig. 2, rather than directly predicting the full attacker trajectory, we first predict a collision pattern, a compact and interpretable representation of the final spatial configuration of the ego and attacker vehicles at the collision moment. The predicted pattern is then used as a target for trajectory realization by a quintic polynomial planner. Three trainable modules, Ego Position Prediction, Attacker Offset Prediction, and Attacker Selection, operate on the encoded scene and condition features to predict the collision pattern.
@@ V Experimental Results
We evaluate our method through comparisons with conditional baselines, ablation studies, and planner-based assessments on the proposed COLLIDE dataset.
@@ V-A Baselines and Evaluation Setup
To enable fair comparisons, we implement three conditional scenario generation baselines, C-SGAN, C-MID, and C-STRIVE, by adopting representative trajectory prediction models, Social-GAN [20], MID [10], and the traffic dynamic model of STRIVE [7], respectively. We concatenate a one-hot vector of collision type and normalized time-to-accident (TTA) as the input of the decoder for the baselines.
To ensure a fair comparison, baselines are extended to the same prediction horizon by applying open-loop autoregression: each iteration predicts 4 frames and appends them to the input sequence for the next step. Our method uses the same history input to directly predict a compact collision pattern, the terminal spatial configuration of the ego and attacker. A quintic polynomial planner then generates the full attacker trajectory to meet the predicted pattern at the desired TTA. This difference in generation strategy (pattern prediction + planner vs. autoregressive rollout) is explicitly accounted for in the evaluation by enforcing the same input horizon and final prediction horizon for all methods.
@@ V-B Evaluation Metrics
Setting
Displacement
Error (m)
Angle
Distance ()
Collision
Rate
Similarity
@@
V-C Results of Conditional Collision Scenario Generation
@@
V-D Ablation Study on Ego Bounding Box Prediction
@@
V-E Planner Evaluation under Safety-Critical Scenarios
@@
Both the IDM and STRIVE rule-based planners rely on the original ego trajectory as a reference path. The IDM planner treats the attacker as a lead vehicle and follows a conservative car-following policy, often resulting in early braking and thus lower collision rates. In contrast, the STRIVE rule-based planner applies a simple brake–accelerate rule to slightly adjust the replayed speed, producing behavior more similar to the original collision scenario. The PDM planner adopts a more complex planning pipeline: it uses the lane centerline as a reference path, generates candidate maneuvers in both lateral and longitudinal directions, and evaluates them using an internal scoring function. To assess potential collisions, it simulates the future motion of surrounding vehicles under the assumption that each continues with its current velocity rather than relying on learned trajectory predictors. This uncertainty in reasoning about the front agent leads to a higher collision rate.
@@
V-F Improving Planner with Generated Scenarios
@@
Results. C-STRIVE tends to yield overly aggressive attacker behaviors, often leading to collisions earlier than the specified TTA. This drives the PDM planner to favor more frequent lane changes, which is advantageous in scenarios such as Opposite Direction (OD) and Rear-End (RE), where collision avoidance is only feasible through lateral maneuvers. These results highlight that different scenario generation methods expose different planner failure modes, underscoring the need for a broader and more controllable set of testing scenarios to comprehensively assess planner robustness and generalization.
TABLE V: Parameter tuning results on planners. We tune the policy parameters of PDM using scenarios generated by different methods. The table reports the collision rate of PDM across five collision types.

### v3

@@ Abstract
Evaluating the safety of autonomous vehicles (AVs) requires diverse, safety-critical scenarios, with collisions being especially important yet rare and unsafe to collect in the real world. Therefore, the community has been focusing on generating safety-critical scenarios in simulation. However, controlling attributes such as collision type and time-to-accident (TTA) remains challenging. We introduce a new task called controllable collision scenario generation, where the goal is to produce trajectories that realize a user-specified collision type and TTA, to investigate the feasibility of automatically generating desired collision scenarios. To support this task, we present COLLIDE, a large-scale collision scenario dataset constructed by transforming real-world driving logs into diverse collisions, balanced across five representative collision types and different TTA intervals. We propose a framework that predicts Collision Pattern, a compact and interpretable representation that captures the spatial configuration of the ego and the adversarial vehicles at impact, before rolling out full adversarial trajectories. Experiments show that our approach outperforms strong baselines in both collision rate and controllability. Furthermore, generated scenarios consistently induce higher planner failure rates, revealing limitations of existing planners. We demonstrate that these scenarios fine-tune planners for robustness improvements, contributing to safer AV deployment in different collision scenarios. Additional generated scenarios are available at this project webpage.
@@ II Related Work
Scenario Generation. Existing approaches can be categorized as data-driven, adversarial, and knowledge-based, according to [1]. Data-driven methods [23, 24, 25] extract rare collision events from large-scale real-world datasets [26], but such events are sparse and highly imbalanced in terms of both collision type and urgency (TTA), limiting their value for training or evaluation. Adversarial methods [7, 27, 8] perturb agent trajectories within learned traffic dynamics models [28] to induce failures, but they often bias toward specific failure modes and lack explicit controllability over collision semantics. Knowledge-based approaches [29, 30, 17, 31] incorporate rule-based priors or pre-defined templates to guarantee coverage of certain collision types but are not scalable, as increasing intra-class variation requires substantial manual effort and does not generalize efficiently.
@@ IV-B Framework Overview
[…]
@@ V-A Baselines and Evaluation Setup
We compare our method with three conditional scenario generation baselines, C-SGAN, C-MID, and C-STRIVE, implemented using Social-GAN [20], MID [10], and the traffic dynamic model of STRIVE [7], respectively. For all baselines, a one-hot vector of collision type and the normalized time-to-accident (TTA) are concatenated as conditional inputs to the decoder. To ensure a fair comparison, baselines are extended to the same prediction horizon by applying open-loop autoregression: each iteration predicts 4 frames and appends them to the input sequence for the next step. Our method uses the same history input to directly predict a compact collision pattern, the terminal spatial configuration of the ego and attacker. A quintic polynomial planner then generates the full attacker trajectory to meet the predicted pattern at the desired TTA. This difference in generation strategy (pattern prediction + planner vs. autoregressive rollout) is explicitly accounted for in the evaluation by enforcing the same input horizon and final prediction horizon for all methods.
@@ V-B Evaluation Metrics
V-C Implementation and Training Details
Three modules, Ego Position Prediction, Attacker Offset Prediction, and Attacker Selection, are trained jointly in an end-to-end manner using supervision from ground-truth collision patterns extracted from COLLIDE. The full training loss is defined as a weighted sum of the ego position prediction loss, attacker offset regression loss, and attacker selection classification loss. We train the network for 1000 epochs using the Adam optimizer with a learning rate of $5\times 10^{-5}$ and a batch size of 128. The overall inference latency is approximately 0.37 ms per scene on a single GPU, including both collision pattern prediction and polynomial trajectory realization.
@@ V-C Implementation and Training Details
Setting Displacement Error (m) Angle Distance () Collision Rate Similarity
@@
V-D Results of Conditional Collision Scenario Generation
@@
V-E Ablation Study on Ego Bounding Box Prediction
@@
V-F Planner Evaluation under Safety-Critical Scenarios
@@
Both the IDM and STRIVE rule-based planners rely on the original ego trajectory as a reference path. The IDM planner treats the attacker as a lead vehicle and follows a conservative car-following policy, often resulting in early braking and thus lower collision rates. In contrast, the STRIVE rule-based planner applies a simple brake–accelerate rule to slightly adjust the replayed speed, producing behavior more similar to the original collision scenario. The PDM planner uses the lane centerline as a reference path, generates candidate maneuvers in both lateral and longitudinal directions, and evaluates them using an internal scoring function. To assess potential collisions, it simulates the future motion of surrounding vehicles under the assumption that each continues with its current velocity rather than relying on learned trajectory predictors. This uncertainty in reasoning about the front agent leads to a higher collision rate.
@@
V-G Improving Planner with Generated Scenarios
@@
Results. C-STRIVE tends to yield overly aggressive attacker behaviors, often leading to collisions earlier than the specified TTA. This drives the PDM planner to favor more frequent lane changes, which is advantageous in scenarios such as Opposite Direction (OD) and Rear-End (RE), where collision avoidance is only feasible through lateral maneuvers. These results highlight that different scenario generation methods expose different planner failure modes, underscoring the need for a broader and more controllable set of testing scenarios to comprehensively assess planner robustness and generalization. Qualitative visualizations of planner behavior after fine-tuning with the generated scenarios are provided in the project webpage.
TABLE V: Parameter tuning results on planners. We tune the policy parameters of PDM to evaluate collision rate using scenarios generated by different methods.
@@ VI Conclusion and Future Works
Acknowledgement
This work is sponsored in part by the National Science and Technology Council under grants 113-2628-E-A49-022- and 114-2628-E-A49-007-, MobileDrive, the Higher Education Sprout Project of National Yang Ming Chiao Tung University, the Ministry of Education, and the Yushan Fellow Program Administrative Support Grant.